OpenStudy (anonymous):

Log7x = 1/3log7 64 + 1/2log7 121. How do you solve for x?

OpenStudy (anonymous):

\[\log_{7}x = \log_{7}64 + 1/2\log_{7}121 \]

OpenStudy (anonymous):

There is a 1/3 infront of log7 64, sorry.

OpenStudy (anonymous):

if 1/3 is there the answer is x=44; if x/3 is not there then ans: x=64*11

OpenStudy (anonymous):

How, though?

OpenStudy (anonymous):

use two formulas: m*log(a)=log(a^m); log(a)+log(b)=log(a*b)

OpenStudy (anonymous):

Thank you, Ryan. (:

OpenStudy (anonymous):

NP